OpenStudy (anonymous):

what is the equation for (8,-2) slope =-3

OpenStudy (anonymous):

y + 2 = -3(x-8) y = -3x + 22

OpenStudy (anonymous):

y = -3x+22 kanwar was right~

OpenStudy (anonymous):

Use point slope form to figure this problem out. (y-y1) = m(x-x1) (y-(-2)) = -3(x-(8)) Distribute. y+2 = -3x+24 Subtract 2 from both sides. y = -3x+22